OpenStudy (angelic):

norm must pay to have trash collected. he can put out 1 barrel of trash per week. a rectangular container measures 32 inches by 18 inches by 10 inches and a cylindrical container measures 30 inches high by 20 inches wide, both are on sale for 9.99 each. which holds the most trash

OpenStudy (anonymous):

rectangle-> width*depth* height cylinder -> ((4/pi)*diameter^2)* height

OpenStudy (anonymous):

Volume of cylinder = pi R^2 * H = pi * 10^2 * 30 = pi * 100 * 30 = 3000 pi = 9425 cubic inches